Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 584
Entry time: 03 Feb 2025, 09:39
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 03.02.2025 |
|
|
General Announcement:
- Experiment tuning week is ongoing; Two downtimes happened over the weekend due to 1. an electronic component part of quench circuit of a cold magnet in A20 (11hs) and 2. interlock issues at the injector
laser (3hs).
CTRL:
- Nothing to report.
DA:
- Nothing to report.
EEE:
- Nothing major to report.
VAC:
- Nothing to report.
XPD:
- Nothing to report;
- Naresh K. work went fine on Saturday (wavefront sensor at HED); program was cut because of short time due to downtime.
DET:
- Nothing to report.
ITDM:
- Nothing to report.
XRO:
- No major problems; hiccups were investigated and there is nothing preventing operation.
LAS:
- Not present.
UNSYS:
- Nothing to report.
SPB:
- Nothing to report.
FXE:
- Nothing to report.
MID:
- Nothing to report.
HED:
- Not present.
SCS:
- Nothing to report.
SQS:
- Not present.
SXP:
- Not present. |